OpenStudy (idealist10):

Find the point(s) on the curve of 4x^2+9y^2=36 at which the curvature is largest.

OpenStudy (idealist10):

\[k(x)=\frac{ \left| y'' \right| }{ [1+(y')^2]^{3/2} }\]

OpenStudy (idealist10):

8x+18y(dy/dx)=0 18y(dy/dx)=-8x dy/dx=-8x/18y=-4x/9y @SithsAndGiggles

OpenStudy (anonymous):

Are you having trouble with the second derivative or something?

OpenStudy (anonymous):

\[4x^2+9y^2=36~~\implies~~y=\pm\frac{1}{3}\sqrt{36-4x^2}\] You can also consider the ellipse as a function of \(y\): \[x=\pm\frac{3}{2}\sqrt{4-y^2}\] and find the curvature \[k(y)=\frac{|x''|}{(1+(x')^2)^{3/2}}\] (which you might have to do anyway, judging by the shape of this ellipse).

OpenStudy (irishboy123):

.